In printing an article of 48,000 words, a printer decides to use two sizes of type. Using the larger type, a printed page contains 1,800 words. Using smaller type, a page contains 2,400 words. The article is allotted 21 full pages in a magazine. How many pages must be in smaller type?
Answer by rmromero(383)

What is asked in the problem?
How many pages must be in smaller type?


Given :
An article of 48, 000 words
the larger type, the printed page contains 1,800 words.
using smaller type the page contains 2,400.
the article is alloted 21 full pages in a magazine.


Representation:
Let L = number of pages printed Larger type
S = number of pages printed smaller type


Translate given sentences to mathematical sentences.
the article is alloted 21 full pages in a magazine. -->L + S = 21
1,800L + 2, 400S = 48, 000, divide 100
18L + 24S = 480


Solve:
L + S = 21
18L + 24S = 480


-18 (L + S = 21)
18L + 24S = 480


-18L - 18S = 378
18L + 24S = 480


6S = 102
S = 17 --> pages printed in smaller type